Find the area bounded bounded by the following curves, y=x2/4 and x+4=2y.


1
Expert's answer
2021-08-02T13:51:14-0400

Solution:

"y_1={x^2\over4}"

"{2\over 2}y_2=x+4 \implies y_2={1\over 2}x+{4\over 2}\implies y_2={1\over 2}x+2"

At point of intersection, "y_1=y_2"

"\implies {x^2\over4}={1\over 2}x+2"

Multiply through by "8"

"\implies {x^2\over4}(8)={1\over 2}(8)x+2(8)"

"\implies 2x^2=4x+16"

Divide through by "2"

"\implies { 2\over 2}x^2={4\over 2}x+{16\over 2}"

"\implies x^2=2x+8 \implies x^2-2x-8 =0"

Solving quadratic equation:

"x^2-2x-8 =0"

"(x-4)(x+2)=0"

"\implies x_1-4=0 \implies x_1=4"

"\implies x_2+2=0 \implies x_2=-2"

"Area=\int_{-2}^0(y_2-y_1)dx+\int_0^4(y_2-y_1)dx"

"=\int_{-2}^0({1\over 2}x+2-{x^2\over4})dx+\int_0^4({1\over 2}x+2-{x^2\over4})dx"

"=[{x^2\over 2\times 2}+2x-{x^3\over 4\times 3}]_{-2}^0+[{x^2\over 2\times 2}+2x-{x^3\over 4\times 3}]_0^4"

"=[{x^2\over 4}+2x-{x^3\over 12}]_{-2}^0+[{x^2\over 4}+2x-{x^3\over 12}]_0^4"

"=({0^2\over 4}+2(0)-{0^3\over 12})-({(-2)^2\over 4}+2(-2)-{(-2)^3\over 12})+({4^2\over 4}+2(4)-{4^3\over 12})-({0^2\over 4}+2(0)-{(0)^3\over 12})"

"=(0+0-0)-(1-4+0.666666666)+(4+8-5.333333333)-(0+0-0)"

"=4-1-0.666666666+4+8-5.333333333"

"=9.000000001" square units